# BioT5+: Towards Generalized Biological Understanding with IUPAC Integration and Multi-task Tuning

### v1

@@ 4.2.3 Protein-related Interaction Prediction
Model B-2normal-$\uparrow$ B-4normal-$\uparrow$ R-1normal-$\uparrow$ R-2normal-$\uparrow$ R-Lnormal-$\uparrow$ METEORnormal-$\uparrow$
BioT5+(single task)
0.671 0.597 0.715 0.590 0.655 0.687
BioT5+(single task) wo IUPAC
0.661 0.584 0.706 0.578 0.647 0.677
BioT5+(single task) wo additional data
0.666 0.591 0.711 0.586 0.651 0.681
BioT5+
0.666 0.591 0.710 0.584 0.650 0.681
[Uncaptioned image] Figure 2: Performance (ROUGE-L) comparison on protein description generation tasks.
4.2.4 Description-guided Protein Design
4.3 Ablation Studies
In this section, we conduct ablation studies to investigate the effectiveness of our design in BioT5+. Specifically, we focus on the following 3 scenarios: (1) Do not incorporate IUPAC name of the molecule. As shown in Table 7, removing the IUPAC name results in a noticeable performance drop in the molecule description generation task. This decline highlights the significant role that IUPAC name plays in tasks related to molecular understanding. (2) Do not add PubMed Central and bioRxiv data in pre-training. Results in Table 7 and Table 12 indicate that these two datasets play a crucial role in enhancing molecular understanding. The omission of them leads to a slight but noticeable decrease in performance on molecule description generation and description-guided molecule design tasks. (3) Use T5 default tokenizer for numbers instead of character-based tokenizer. The results in Table 9 demonstrate that the character-based approach for tokenizing numbers is more effective than the default T5 tokenizer on regression tasks.
@@ 7 Ethical Considerations
While BioT5+ presents a significant advancement in bioinformatics and computational biology, its capabilities, particularly in generating molecules based on textual descriptions and predicting chemical reaction products, raise important ethical considerations. One of the primary concerns is the potential misuse of this technology to generate harmful or dangerous molecules, which could pose risks to public safety and environmental health. Moreover, the ability of BioT5+ to predict and generate novel molecular structures may also lead to issues surrounding intellectual property rights and patenting. The ease with which new compounds can be designed and synthesized using AI-driven methods could potentially disrupt traditional research practices and raise questions about the ownership of these discoveries.
@@ Appendix A Additional Results
Model BLEU-2normal-$\uparrow$ BLEU-4normal-$\downarrow$ ROUGE-1normal-$\uparrow$ ROUGE-2normal-$\uparrow$ ROUGE-Lnormal-$\uparrow$ METEORnormal-$\uparrow$
Molecule Description Generation
Alpaca 0.068 0.014 0.178 0.041 0.136 0.107
Baize 0.064 0.015 0.189 0.053 0.148 0.106
ChatGLM 0.055 0.011 0.163 0.036 0.121 0.105
LLaMa 0.059 0.014 0.164 0.066 0.148 0.184
Vicuna 0.052 0.011 0.151 0.055 0.130 0.168
Galactica 0.024 0.008 0.074 0.015 0.063 0.065
Mol-Instructions 0.217 0.143 0.337 0.196 0.291 0.254
Text+Chem T5 0.062 0.036 0.126 0.075 0.119 0.139
MolT5 0.002 0.001 0.036 0.001 0.034 0.033
BioT5+ 0.549 0.497 0.758 0.701 0.747 0.715
Model Exactnormal-$\uparrow$ BLEUnormal-$\uparrow$ Levenshteinnormal-$\downarrow$ RDK FTSnormal-$\uparrow$ MACCS FTSnormal-$\uparrow$ Morgan FTSnormal-$\uparrow$ Validitynormal-$\uparrow$
Description-guided Molecule Design
Alpaca 0.000 0.004 51.088 0.006 0.029 0.000 0.002
Baize 0.000 0.006 53.796 0.000 0.000 0.000 0.002
ChatGLM 0.000 0.004 53.157 0.005 0.000 0.000 0.005
Llama 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Galactica 0.000 0.192 44.152 0.135 0.248 0.088 0.992
Mol-Instructions 0.002 0.345 41.367 0.231 0.412 0.147 1.000
Text+Chem T5 0.097 0.508 41.819 0.352 0.474 0.353 0.721
MolT5 0.112 0.546 38.276 0.400 0.538 0.295 0.773
BioT5+ 0.079 0.795 30.728 0.567 0.687 0.410 1.000
@@ Appendix B Hyper-parameters
The pre-training process spans $300$K steps and is executed on eight NVIDIA 80GB A100 GPUs with batch size $96$ per GPU. The dropout rate is maintained at $0.0$ and the maximum input length during pre-training is set at $512$. Optimization is performed using the AdamW (Loshchilov and Hutter, 2019) optimizer with Root Mean Square scaling. A cosine annealing learning rate scheduler is employed, with the base rate set at $21e-$ 2 and the minimum rate at $51e-$ 5. For multi-task fine-tuning, the dropout rate is searched in [0.0, 0.05, 0.1], and the learning rate is searched in [5e-5, 1e-4, 2e-4, 5e-4]. The total number of steps is $100$K and warmup steps is 6% of total steps. The batch size is set to $768$ for molecule-oriented tasks and $96$ for protein-oriented tasks. The best hyper-parameters for molecule-oriented and protein-oriented task are shown in Table 15.
@@ Appendix D IUPAC Incorporation
[…]
@@ D.1 Mapping Process
Appendix E Fine-tuning Details
@@
E.1 Molecule-oriented Tasks
E.1.1 Molecule Property Prediction
@@
We compare our BioT5+ with LLM-based generalist models, including Alpaca Taori et al. (2023), Baize Xu et al. (2023a), Llama2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), Mol-Instructions Fang et al. (2023), and InstructMol Cao et al. (2023). The baseline results come from InstructMol Cao et al. (2023). The evaluation metric is MAE.
E.1.2 Chemical Reaction-related Tasks
@@
E.1.3 Molecule Description Generation
@@
E.1.4 Description-guided Molecule Design
@@
The compared baselines are a subset of that listed in Section E.1.3, with results mainly sourced from BioT5 Pei et al. (2023). The evaluation metrics are BLEU Papineni et al. (2002), exact match, Levenshtein distance (Miller et al., 2009), three molecular fingerprints (FTS) similarity scores including MACCS (Durant et al., 2002), RDK (Schneider et al., 2015), and Morgan (Rogers and Hahn, 2010), FCD score (Preuer et al., 2018), Text2Mol score Edwards et al. (2021), and validity score.
E.2 Protein-oriented Tasks
E.2.1 Protein Description Generation
@@
E.3 Description-guided Protein Design
@@
E.4 Protein Property Prediction
@@
E.5 Protein-Protein Interaction
@@
The compared baselines are the same with that in Section E.4, with results derived from PEER Xu et al. (2022). The evaluation metric is Accuracy.
E.6 Drug-Target Interaction
@@
Appendix F Case Studies
F.1 Molecule-oriented Cases
@@
Target
The molecule is an L-alpha-amino acid zwitterion arising from transfer of a proton from the carboxy to the amino group of O-ureido-L-serine. It is a tautomer of an O-ureido-L-serine.
@@
Galactica
The molecule is a zwitterion, so it has a positive charge on the nitrogen atom and a negative charge on the oxygen atom. The positive charge is on the nitrogen atom and the negative charge is on the oxygen atom.
Mol-Instructions
The molecule is an amino acid zwitterion obtained by transfer of a proton from the carboxy to the amino group of N-methyl-L-aspartic acid. It is an amino acid zwitterion and a N-methyl-L-aspartate. It is functionally related to a N-methyl-L-aspartic acid. It is a tautomer of a N-methyl-L-aspartic acid. It is a tautomer of a N-methyl-L-aspartic acid zw
Text+Chem T5
The molecule is an N-acyl-D-alpha-amino acid zwitterion obtained by transfer of a proton from the carboxy to the amino group of N-carbamoyl-D-ornithine; major species at pH 7.3. It is a tautomer of a N-carbamoyl-D-ornithine.
BioT5+
The molecule is an amino acid zwitterion resulting from the transfer of a proton from the carboxy to the amino group of O-ureido-L-serine. It is a tautomer of an O-ureido-L-serine.
Input
[O][=C][Branch1][C][O][C@H1][Branch1][C][O][C][=Branch1]
@@
Target
The molecule is the ketoaldonic acid phosphate formed formally from D-erythronic acid by oxidation of the 3 -hydroxy group to an oxo group and phosphorylation at the 1 -hydroxy group. It is a conjugate acid of a (R)-2 -hydroxy-3 -oxo-4 -(phosphonatooxy)butanoate(3 -). It is an enantiomer of a (S)-2 -hydroxy-3 -oxo-4 -(phosphonooxy)butanoic acid.
@@
Galactica
The molecule is a carboxylic acid.
Mol-Instructions
The molecule is a metabolite found in Escherichia coli. It has a role as a human metabolite and a Saccharomyces cerevisiae metabolite. It is a conjugate acid of a 3-phosphoglycerate(2-). It is an enantiomer of a 3-phosphoglycerate(1-). It is a tautomer of a 3-phosphoglyceric acid zwitterion. It is a tautomer of a 3-phosphoglyceric acid zwitterion.
Text+Chem T5
The molecule is the D-enantiomer of 2-dehydro-3-deoxy-D-glucuronic acid. It has a role as an Escherichia coli metabolite and a mouse metabolite. It is a conjugate acid of a (R)-2-dehydro-3-deoxy-D-glucuronate(2-). It is an enantiomer of a (S)-2-dehydro-3-deoxy-L-glucuronic acid.
BioT5+
The molecule is the ketoaldonic acid phosphate formed formally from L-erythronic acid by oxidation of the 3 -hydroxy group to an oxo group and phosphorylation at the 1 -hydroxy group. It derives from a L-erythronic acid. It is a conjugate acid of a (R)-2 -hydroxy-3 -oxo-4 -(phosphonatooxy)butanoate(3 -). It is an enantiomer of a (S)-2 -hydroxy-3 -oxo-4 -(phosphonooxy)butanoic acid.
Input
[NH3+1][C@@H1][Branch1][#C][C][C][=C][NH1][C][=C][C][=C][C]
@@
Target
The molecule is the L-enantiomer of tryptophanium. It has a role as an animal metabolite and a plant metabolite. It is a conjugate acid of a L-tryptophan. It is an enantiomer of a D-tryptophanium.
@@
Galactica
The molecule is [NH3+1][C@@H1](CC1=C[NH1]C2=CC=CC=C12)C(=O)O.
Mol-Instructions
The molecule is an amino acid zwitterion. It has a role as a human metabolite. It is functionally related to an alanine. It is a tautomer of a L-alanine. It contains a L-alanine. It is a tautomer of a D-alanine zwitterion. It is an enantiomer of a D-alanine zwitterion. It is a tautomer of a D-alanine zwitterion. It is an enantiomer of a L-alanine zwitterion.
Text+Chem T5 (Ours)
The molecule is an ammonium ion resulting from the protonation of the amino group of (S)-nefopam. It is a conjugate acid of a (S)-nefopam. It is an enantiomer of a (R)-nefopam(1+).
BioT5+
The molecule is the L-enantiomer of tryptophanium. It has a role as an Escherichia coli metabolite and a Saccharomyces cerevisiae metabolite. It is a conjugate acid of a L-tryptophan. It is an enantiomer of a D-tryptophanium.
@@
Input
The molecule is any secondary alcohol that is one of the eight possible diastereoisomers of 5 -methyl-2 -(propan-2 -yl)cyclohexan-1 -ol. It has a role as a volatile oil component. It is a p-menthane monoterpenoid and a secondary alcohol.
Target SMILES
CC1CCC(C(C)C)C(O)C1
Target SELFIES
[C][C][C][C][C][Branch1][=Branch1][C][Branch1]
@@
Galactica
CC1=CC=C(C)C(C)=C1NC(=O)C1=CC=C(Cl)C=C1
Mol-Instructions
[C][C@@H1][C][C][C@@H1][Branch1]
Text+Chem T5
CC(C)C1CCC(CC1)O
BioT5+
[C][C][C][C][C][Branch1][=Branch1][C][Branch1]
@@
Input
The molecule is a phenylsulfate oxoanion that is the conjugate base of 2 -aminophenyl hydrogen sulfate, obtained by deprotonation of the sulfo group; major species at pH 7.3. It is a conjugate base of a 2 -aminophenyl hydrogen sulfate.
Target SMILES
NC1=CC=CC=C1OS(=O)(=O)[O-1]
Target SELFIES
[N][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Mol-Instructions
[C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1]
@@
Text+Chem T5
C1=CC=C(C(=C1)N)OS(=O)(=O)[O-]
BioT5+
[N][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Input
The molecule is a dicarboxylic acid. It is a conjugate acid of a diphenate(1 -). It derives from a hydride of a biphenyl.
Target SMILES
O=C(O)C1=CC=CC=C1C2=CC=CC=C2C(=O)O
Target SELFIES
[O][=C][Branch1][C][O][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Galactica
C[C@@H]1[C@@H]2C[C@H]1[C@@H]1O[C@@H]12
Mol-Instructions
[C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][C][=Branch1]
@@
Text+Chem T5
C1=CC=C(C(=C1)C2=CC=CC=C2)C(=O)O
BioT5+
[O][=C][Branch1][C][O][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Input
[O][=N+1][Branch1][C][O-1][O-1].[O][=C][NH1][C][=Branch1][C][=O][C][=C][C][=C][C][=C]
@@
Target SMILES
O=S(=O)(O)O.[K+]
Target SELFIES
[O][=S][=Branch1][C][=O][Branch1][C][O][O].[K+1]
@@
Galactica
O=C1[NH1]C(=O)C2=CC=CC=C2[NH1]1
Mol-Instructions
[C][C][=Branch1][C][=O][O].[C][C][=Branch1][C][=O][O].[C][C][=Branch1][C][=O][O].[O]
@@
BioT5+
[O][=S][=Branch1][C][=O][Branch1][C][O][O].[K+1]

### v2

@@ 4.1.2 Chemical Reaction-related Tasks
[…]
BioT5+(single task) 0.671 0.597 0.715 0.590 0.655 0.687
BioT5+(single task) wo IUPAC 0.661 0.584 0.706 0.578 0.647 0.677
BioT5+(single task) wo additional data 0.666 0.591 0.711 0.586 0.651 0.681
BioT5+ 0.666 0.591 0.710 0.584 0.650 0.681
Figure 2: Performance (ROUGE-L) comparison on protein description generation tasks.
@@
4.1.3 Molecule Description Generation
4.1.4 Description-guided Molecule Design
4.2 Protein-oriented Tasks
4.2.1 Protein Description Generation
4.2.2 Protein Property Prediction
4.2.3 Protein-related Interaction Prediction
4.2.4 Description-guided Protein Design
4.3 Ablation Studies
In this section, we conduct ablation studies to investigate the effectiveness of our design in BioT5+. Specifically, we focus on the following 3 scenarios: (1) Do not incorporate IUPAC name of the molecule. As shown in Table 7, removing the IUPAC name results in a noticeable performance drop in the molecule description generation task. This decline highlights the significant role that IUPAC name plays in tasks related to molecular understanding. (2) Do not add PubMed Central and bioRxiv data in pre-training. Results in Table 7 and Table 13 indicate that these two datasets play a crucial role in enhancing molecular understanding. The omission of them leads to a slight but noticeable decrease in performance on molecule description generation and description-guided molecule design tasks. (3) Use T5 default tokenizer for numbers instead of character-based tokenizer. The results in Table 9 demonstrate that the character-based approach for tokenizing numbers is more effective than the default T5 tokenizer on regression tasks. We also conduct an ablation study to further contrast single-task and multi-task tuning strategies in Appendix Section F.
@@ 7 Ethical Considerations
While BioT5+ presents a significant advancement, its capabilities, particularly in generating molecules based on textual descriptions and predicting chemical reaction products, raise important ethical considerations. One of the concerns is the potential misuse of this technology to generate harmful or dangerous molecules, which could pose risks to public safety and environmental health. Moreover, the ability of BioT5+ to predict and generate novel molecules may also lead to issues surrounding intellectual property rights and patenting. The ease with which new compounds can be designed and synthesized using AI-driven methods could potentially disrupt traditional research practices and raise questions about the ownership of these discoveries.
8 Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the "Double-First Class" Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. Qizhi Pei is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China.
@@ Appendix B Hyper-parameters
Pre-training.
The pre-training process spans $300$K steps and is executed on eight NVIDIA 80GB A100 GPUs with batch size $96$ per GPU. To balance the data from different tasks during pre-training, we adopt a batch-level balancing strategy, where each batch evenly includes data from all eight different tasks, ensuring a more balanced and comprehensive pre-training process. For small datasets, such as molecule-text pairs and protein-text pairs, we employ a round-robin strategy to repeat their usage multiple times, compensating for their limited size. The dropout rate is maintained at $0.0$ and the maximum input length during pre-training is set at $512$. Optimization is performed using the AdamW (Loshchilov and Hutter, 2019) optimizer with Root Mean Square scaling. A cosine annealing learning rate scheduler is employed, with the base rate set at $21e-$ 2 and the minimum rate at $51e-$ 5.
Multi-task Fine-tuning.
For multi-task fine-tuning, the dropout rate is searched in [0.0, 0.05, 0.1], and the learning rate is searched in [5e-5, 1e-4, 2e-4, 5e-4]. The total number of steps is $100$K and warmup steps is 6% of total steps. The batch size is set to $768$ for molecule-oriented tasks and $96$ for protein-oriented tasks. The best hyper-parameters for molecule-oriented and protein-oriented task are shown in Table 12.
@@
Table 13: Ablation of additional data on the description-guided molecule design task.
[…]
BioT5+(single task) 0.877 0.535 12.777 0.909 0.842 0.784 0.350 0.580 1.000
BioT5+(single task) wo additional data 0.875 0.516 12.840 0.904 0.833 0.777 0.358 0.579 1.000
BioT5+ 0.872 0.522 12.776 0.907 0.835 0.779 0.353 0.579 1.000
Table 14: Performance comparison of single-task and multi-task versions of BioT5+ across various datasets and tasks. Mol2text means molecule description generation, and text2mol means description-guided molecule design.
Dataset/Task Task Type Single-task Multi-task Metric
BACE classification 87.8 86.2 AUROC $\uparrow$
BBBP classification 78.5 76.5 AUROC $\uparrow$
Clintox classification 97.0 92.3 AUROC $\uparrow$
ChEBI-20-mol2text generation 0.687 0.681 METEOR $\uparrow$
ChEBI-20-text2mol generation 0.877/0.535 0.872/0.522 BLEU/Exact Match $\uparrow$
HOMO regression 0.0029 0.0022 MAE $\downarrow$
LUMO regression 0.0029 0.0024 MAE $\downarrow$
HOMO-LUMO gap regression 0.0040 0.0028 MAE $\downarrow$
[…]
Molecule Description Generation
Alpaca-7B 0.068 0.014 0.178 0.041 0.136 0.107
Baize-7B 0.064 0.015 0.189 0.053 0.148 0.106
ChatGLM-6B 0.055 0.011 0.163 0.036 0.121 0.105
Llama-7B 0.059 0.014 0.164 0.066 0.148 0.184
Vicuna-7B 0.052 0.011 0.151 0.055 0.130 0.168
Galactica-6.7B 0.024 0.008 0.074 0.015 0.063 0.065
Mol-Instructions-7B 0.217 0.143 0.337 0.196 0.291 0.254
Text+Chem T5-223M 0.062 0.036 0.126 0.075 0.119 0.139
MolT5-248M 0.002 0.001 0.036 0.001 0.034 0.033
BioT5+ 0.549 0.497 0.758 0.701 0.747 0.715
[…]
Description-guided Molecule Design
Alpaca-7B 0.000 0.004 51.088 0.006 0.029 0.000 0.002
Baize-7B 0.000 0.006 53.796 0.000 0.000 0.000 0.002
ChatGLM-6B 0.000 0.004 53.157 0.005 0.000 0.000 0.005
Llama-7B 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna-7B 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Galactica-6.7B 0.000 0.192 44.152 0.135 0.248 0.088 0.992
Mol-Instructions-7B 0.002 0.345 41.367 0.231 0.412 0.147 1.000
Text+Chem T5-223M 0.097 0.508 41.819 0.352 0.474 0.353 0.721
MolT5-248M 0.112 0.546 38.276 0.400 0.538 0.295 0.773
BioT5+ 0.079 0.795 30.728 0.567 0.687 0.410 1.000
@@ Appendix D IUPAC Incorporation
In our downstream tasks of molecule property prediction and molecule description generation, it is effective to enrich molecules with their IUPAC name. One reason is that IUPAC names are more commonly found in bio-text. By explicitly incorporating IUPAC names in the molecular context during pre-training, the model more readily learns relevant molecular knowledge and establishes connections between the molecule and its contextual information. Additionally, IUPAC names inherently contain structural information about the molecule, such as functional groups and other structural components. This information allows the model to better understand the molecular structure, thus predicting molecule properties with higher accuracy and generating more accurate and detailed descriptions.
@@ D.1 Mapping Process
Appendix E Comparison with BioT5
Regarding the size of pre-training data and preprocessing techniques, BioT5+ introduces several enhancements over BioT5 Pei et al. (2023):
  • Addition of 28.8M PubChem molecular data, including molecule SELFIES along with their IUPAC names.
  • Addition of 28.8M full articles from PubMed.
  • Addition of 2.3M bioRxiv abstracts as wrapped biotext.
  • In the molecule-text bidirectional translation task, IUPAC representations are included in the text description.
  • For wrapped biotext pre-training, detected molecule names in BioT5 are directly replaced with corresponding SELFIES. In BioT5+, we first determine whether a molecule name is an IUPAC name; if so, it is appended with the corresponding SELFIES. If not, both the IUPAC name and SELFIES are appended. Besides, we further ensure data quality by only processing detected bioentities with a confidence score (from BERN2 Sung et al. (2022)) above 0.9.
  • The numbers that appear in the pre-training corpus are tokenized character by character in BioT5+’s tokenizer.
Appendix F Additional Ablation Study
To contrast single-task and multi-task tuning strategies, we further fine-tune BioT5+ with a single-task setting on three MoleculeNet Wu et al. (2018) classification tasks including BACE, BBBP, Clintox, and three QM9 Fang et al. (2023) regression tasks including HOMO, LUMO, and HOMO-LUMO gap. The consolidated results are shown in Table 14 (we also summarize the results in Table 7 and Table 13 for generation tasks on ChEBI-20 Edwards et al. (2022) dataset here), covering 3 types of tasks: MoleculeNet for classification task, ChEBI-20 for generation task, and QM9 for regression task. Our findings indicate that: (1) Single-task fine-tuning yields different performance in different tasks. For the tasks such as BACE, BBBP, Clintox, and ChEBI-20, single-task tuned BioT5+ performs closely or slightly better than the multi-task-tuned version. (2) Multi-task fine-tuning can still be advantageous for tasks with inherent correlations, such as the prediction of HOMO, LUMO, and the HOMO-LUMO gap. This evidence points to the potential for cross-task generalization, even if it is not uniformly applicable across all tasks.
Appendix G Fine-tuning Details
@@
G.1 Molecule-oriented Tasks
G.1.1 Molecule Property Prediction
@@
We compare our BioT5+ with LLM-based generalist models, including Llama2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), Mol-Instructions Fang et al. (2023), and InstructMol Cao et al. (2023). The baseline results come from InstructMol Cao et al. (2023). The evaluation metric is MAE.
G.1.2 Chemical Reaction-related Tasks
@@
G.1.3 Molecule Description Generation
@@
G.1.4 Description-guided Molecule Design
@@
The compared baselines are a subset of that listed in Section G.1.3 except for Llama2 Touvron et al. (2023b), with results mainly sourced from BioT5 Pei et al. (2023) and MolReGPT Li et al. (2024). The evaluation metrics are BLEU Papineni et al. (2002), exact match, Levenshtein distance (Miller et al., 2009), three molecular fingerprints (FTS) similarity scores including MACCS (Durant et al., 2002), RDK (Schneider et al., 2015), and Morgan (Rogers and Hahn, 2010), FCD score (Preuer et al., 2018), Text2Mol score Edwards et al. (2021), and validity score.
G.2 Protein-oriented Tasks
G.2.1 Protein Description Generation
@@
G.3 Description-guided Protein Design
@@
G.4 Protein Property Prediction
@@
G.5 Protein-Protein Interaction
@@
The compared baselines are the same with that in Section G.4, with results derived from PEER Xu et al. (2022). The evaluation metric is Accuracy.
G.6 Drug-Target Interaction
@@
Appendix H Case Studies
In this section we show some cases for our fine-tuned BioT5+ model. We only showcase key parts of inputs, intentionally omitting the instruction context due to the varying requirements for instruction context across different baseline methods.
H.1 Molecule-oriented Cases
@@
Target The molecule is an L-alpha-amino acid zwitterion arising from transfer of a proton from the carboxy to the amino group of O-ureido-L-serine. It is a tautomer of an O-ureido-L-serine.
@@
Galactica The molecule is a zwitterion, so it has a positive charge on the nitrogen atom and a negative charge on the oxygen atom. The positive charge is on the nitrogen atom and the negative charge is on the oxygen atom.
Mol-Instructions The molecule is an amino acid zwitterion obtained by transfer of a proton from the carboxy to the amino group of N-methyl-L-aspartic acid. It is an amino acid zwitterion and a N-methyl-L-aspartate. It is functionally related to a N-methyl-L-aspartic acid. It is a tautomer of a N-methyl-L-aspartic acid. It is a tautomer of a N-methyl-L-aspartic acid zw
Text+Chem T5 The molecule is an N-acyl-D-alpha-amino acid zwitterion obtained by transfer of a proton from the carboxy to the amino group of N-carbamoyl-D-ornithine; major species at pH 7.3. It is a tautomer of a N-carbamoyl-D-ornithine.
BioT5+ The molecule is an amino acid zwitterion resulting from the transfer of a proton from the carboxy to the amino group of O-ureido-L-serine. It is a tautomer of an O-ureido-L-serine.
Input [O][=C][Branch1][C][O][C@H1][Branch1][C][O][C][=Branch1]
@@
Target The molecule is the ketoaldonic acid phosphate formed formally from D-erythronic acid by oxidation of the 3 -hydroxy group to an oxo group and phosphorylation at the 1 -hydroxy group. It is a conjugate acid of a (R)-2 -hydroxy-3 -oxo-4 -(phosphonatooxy)butanoate(3 -). It is an enantiomer of a (S)-2 -hydroxy-3 -oxo-4 -(phosphonooxy)butanoic acid.
@@
Galactica The molecule is a carboxylic acid.
Mol-Instructions The molecule is a metabolite found in Escherichia coli. It has a role as a human metabolite and a Saccharomyces cerevisiae metabolite. It is a conjugate acid of a 3-phosphoglycerate(2-). It is an enantiomer of a 3-phosphoglycerate(1-). It is a tautomer of a 3-phosphoglyceric acid zwitterion. It is a tautomer of a 3-phosphoglyceric acid zwitterion.
Text+Chem T5 The molecule is the D-enantiomer of 2-dehydro-3-deoxy-D-glucuronic acid. It has a role as an Escherichia coli metabolite and a mouse metabolite. It is a conjugate acid of a (R)-2-dehydro-3-deoxy-D-glucuronate(2-). It is an enantiomer of a (S)-2-dehydro-3-deoxy-L-glucuronic acid.
BioT5+ The molecule is the ketoaldonic acid phosphate formed formally from L-erythronic acid by oxidation of the 3 -hydroxy group to an oxo group and phosphorylation at the 1 -hydroxy group. It derives from a L-erythronic acid. It is a conjugate acid of a (R)-2 -hydroxy-3 -oxo-4 -(phosphonatooxy)butanoate(3 -). It is an enantiomer of a (S)-2 -hydroxy-3 -oxo-4 -(phosphonooxy)butanoic acid.
Input [NH3+1][C@@H1][Branch1][#C][C][C][=C][NH1][C][=C][C][=C][C]
@@
Target The molecule is the L-enantiomer of tryptophanium. It has a role as an animal metabolite and a plant metabolite. It is a conjugate acid of a L-tryptophan. It is an enantiomer of a D-tryptophanium.
@@
Galactica The molecule is [NH3+1][C@@H1](CC1=C[NH1]C2=CC=CC=C12)C(=O)O.
Mol-Instructions The molecule is an amino acid zwitterion. It has a role as a human metabolite. It is functionally related to an alanine. It is a tautomer of a L-alanine. It contains a L-alanine. It is a tautomer of a D-alanine zwitterion. It is an enantiomer of a D-alanine zwitterion. It is a tautomer of a D-alanine zwitterion. It is an enantiomer of a L-alanine zwitterion.
Text+Chem T5 (Ours) The molecule is an ammonium ion resulting from the protonation of the amino group of (S)-nefopam. It is a conjugate acid of a (S)-nefopam. It is an enantiomer of a (R)-nefopam(1+).
BioT5+ The molecule is the L-enantiomer of tryptophanium. It has a role as an Escherichia coli metabolite and a Saccharomyces cerevisiae metabolite. It is a conjugate acid of a L-tryptophan. It is an enantiomer of a D-tryptophanium.
@@
Input The molecule is any secondary alcohol that is one of the eight possible diastereoisomers of 5 -methyl-2 -(propan-2 -yl)cyclohexan-1 -ol. It has a role as a volatile oil component. It is a p-menthane monoterpenoid and a secondary alcohol.
Target SMILES CC1CCC(C(C)C)C(O)C1
Target SELFIES [C][C][C][C][C][Branch1][=Branch1][C][Branch1]
@@
Galactica CC1=CC=C(C)C(C)=C1NC(=O)C1=CC=C(Cl)C=C1
Mol-Instructions [C][C@@H1][C][C][C@@H1][Branch1]
Text+Chem T5 CC(C)C1CCC(CC1)O
BioT5+ [C][C][C][C][C][Branch1][=Branch1][C][Branch1]
@@
Input The molecule is a phenylsulfate oxoanion that is the conjugate base of 2 -aminophenyl hydrogen sulfate, obtained by deprotonation of the sulfo group; major species at pH 7.3. It is a conjugate base of a 2 -aminophenyl hydrogen sulfate.
Target SMILES NC1=CC=CC=C1OS(=O)(=O)[O-1]
Target SELFIES [N][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Mol-Instructions [C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1]
@@
Text+Chem T5 C1=CC=C(C(=C1)N)OS(=O)(=O)[O-]
BioT5+ [N][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Input The molecule is a dicarboxylic acid. It is a conjugate acid of a diphenate(1 -). It derives from a hydride of a biphenyl.
Target SMILES O=C(O)C1=CC=CC=C1C2=CC=CC=C2C(=O)O
Target SELFIES [O][=C][Branch1][C][O][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Galactica C[C@@H]1[C@@H]2C[C@H]1[C@@H]1O[C@@H]12
Mol-Instructions [C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][C][=Branch1]
@@
Text+Chem T5 C1=CC=C(C(=C1)C2=CC=CC=C2)C(=O)O
BioT5+ [O][=C][Branch1][C][O][C][=C][C][=C][C][=C][Ring1][=Branch1]
@@
Input [O][=N+1][Branch1][C][O-1][O-1].[O][=C][NH1][C][=Branch1][C][=O][C][=C][C][=C][C][=C]
@@
Target SMILES O=S(=O)(O)O.[K+]
Target SELFIES [O][=S][=Branch1][C][=O][Branch1][C][O][O].[K+1]
@@
Galactica O=C1[NH1]C(=O)C2=CC=CC=C2[NH1]1
Mol-Instructions [C][C][=Branch1][C][=O][O].[C][C][=Branch1][C][=O][O].[C][C][=Branch1][C][=O][O].[O]
@@
BioT5+ [O][=S][=Branch1][C][=O][Branch1][C][O][O].[K+1]